Difference of a number multiplied by 4 and 11 is -3
Answer by wuwei96815(245)

First, we have to translate the language into a math problem. The language is vague, but let's try:

4x -11x = -3
-7x = -3
x = -3/-7 = 3/7

The answer is correct, but only if we interpreted the word problem corrctly. However, it is my opinion that the word problem is vague and can easily be misintrepreted.